# CMB Constraints on Natural Inflation with Gauge Field Production

### v1

@@ 5 Discussions and conclusion
The predictions of the CMB spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ for the axion model are under tension with current observational data - see the continuous line in the Fig. 5. When both Planck18 and BK18 data are used, the model is ruled out at $2$-$\sigma$, whereas it is consistent at $2$-$\sigma$ for $f\gtrsim 5.5$ when only Planck18 data is considered. However, the presence of the gauge field affects the CMB scales, leading to significant impacts on the predictions of the spectral index ($n_{s}$) and the tensor-to-scalar ratio (r) at the CMB scales. As we increase the value of the coupling constant $\alpha$, the spectral index starts to decrease first, but with further increases of the $\alpha$, the spectral index starts to reverse its trend. This seems to be the common feature for all values of $f$. Therefore, for all super-Planckian values of $f$, the model becomes consistent for a certain range of $\alpha$ only. Interestingly enough, the constraints coming from $n_{s}$-$r$ plane are more stringent than the constraints originating from the non-gaussianities or running of the spectrum. In comparison to the previous analysis where an upper limit was derived, we find a range of $\xi_{*}$, or $\alpha$ for which the model is consistent. The existence of both lower and upper limits originates due to the improved limit of $n_{s}$ from the latest data. The limits are summarised in Table  1. In short, a coupling between the gauge field and the pseudoscalar inflaton helps to alleviate the tension of the natural inflation model and makes it consistent with the current observational data.
@@ Acknowledgement
The authors would like to thank Anish Ghoshal for several fruitful discussions, and for taking part in the initial stages of the project. N.J thanks Shu-Lin Cheng for his help in numerical simulations. N.J. was supported by the National Postdoctoral Fellowship of the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India, File No. PDF/2021/004114. K.D would like to acknowledge support from the ICTP, Trieste through the Associates Programme (2023-2028).

### v2

@@ 5 Discussions and conclusion
The predictions of the CMB spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ for the axion model are under tension with current observational data - see the continuous line in the Fig. 5. When both Planck18 and BK18 data are used, the model is ruled out at $2$-$\sigma$, whereas it is consistent at $2$-$\sigma$ for $f\gtrsim 5.5$ when only Planck18 data is considered. However, the presence of the gauge field affects the CMB scales, leading to significant impacts on the predictions of the spectral index ($n_{s}$) and the tensor-to-scalar ratio (r) at the CMB scales. As we increase the value of the coupling constant $\alpha$, the spectral index starts to decrease first, but with further increases of the $\alpha$, the spectral index starts to reverse its trend. This seems to be the common feature for all values of $f$. Therefore, for all super-Planckian values of $f$, the model becomes consistent (in reference to different data combinations) for a certain range of $\alpha$ only. Natural inflation model with no gauge field is excluded at $2$-$\sigma$ for all values of $f$ when viewed w.r.t Planck18+BK18 data. But, for $f=4$, the model becomes consistent for a range of values of $\alpha$ for Planck18+BK18 data. Interestingly enough, the constraints coming from $n_{s}$-$r$ plane are more stringent than the constraints originating from the non-gaussianities or running of the spectrum. In comparison to the previous analysis where an upper limit was derived, we find a range of $\xi_{*}$, or $\alpha$ for which the model is consistent. The existence of both lower and upper limits originates due to the improved limit of $n_{s}$ from the latest data. The limits are summarised in Table  1. In short, a coupling between the gauge field and the pseudoscalar inflaton helps to alleviate the tension of the natural inflation model and makes it consistent with the current observational data.
@@ Acknowledgement
The authors would like to thank Anish Ghoshal for several fruitful discussions, and for taking part in the initial stages of the project. N.J thanks Shu-Lin Cheng for his help in numerical simulations. N.J. was supported by the National Postdoctoral Fellowship of the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India, File No. PDF/2021/004114. K.D would like to acknowledge support from the ICTP, Trieste through the Associates Programme (2023-2028). We thank the anonymous referee for her/his critical comments about the importance of the results, and accordingly, we have presented our results slightly differently.